OpenStudy (anonymous):

what is the rectangular equation equivalent to the polar equation r=sec(theta)? A. x-y=1 B. x=1 C. x+y=1 D. xy=1

OpenStudy (anonymous):

i know that x=rcos(theta) y=rsin(theta) and x^2 +y ^2=r^2

OpenStudy (jdoe0001):

keep in mind that sec = 1/cos

OpenStudy (anonymous):

sorry got it